Write the equation of a parabola with a vertex of (-2,-3) and contains the point (1,2)
Answer by stanbon(75887)

Write the equation of a parabola with a vertex of (-2,-3) and contains the point (1,2)
-----
Assuming the parabola opens up you get:
y = a(x+2)-3
----
Solve for "a":
2 = a(1+2)-3
------
2 = 3a - 3
3a = 5
a = (5/3)
------------------
y = (5/3)(x+2)-3
----
y = (5/3)x + (10/3)-(9/3)
----
y = (5/3)x + (1/3)
=======================
Cheers,
Stan H.
=======================